OpenStudy (anonymous):

which ratios form a portion a.4/5,20/25 b.8/12,18/24 c.1/3,7/24

OpenStudy (er.mohd.amir):

try to simplify each option by diveding

OpenStudy (anonymous):

Diveding what number

OpenStudy (er.mohd.amir):

4/5,20/25 so 20/25=4*5/5*5=4/5

OpenStudy (anonymous):

So is the answer a

OpenStudy (er.mohd.amir):

first solve b also and c

OpenStudy (er.mohd.amir):

yes A is correct ans